OpenStudy (anonymous):

How to do this? As the limit of x approaches 0 ((1/Sqt 1+s) -1) / s

hartnn (hartnn):

do you know how to rationalize ?

hartnn (hartnn):

by multiplying with the conjugate ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but nothing cancels out you still have 0/0

hartnn (hartnn):

ok, so, first step was to cross-multiply in numerator 1/sqrt(s+1) -1 = (1-sqrt(s+1)) / sqrt(s+1) got this ?

OpenStudy (anonymous):

why do you cross multiply..?

hartnn (hartnn):

just a simplification technique....i mean, what else we can do :P

hartnn (hartnn):

now we rationalize the numerator! by multiplying and dividing by 1+sqrt(s+1)

OpenStudy (anonymous):

yes i got that

hartnn (hartnn):

good, so what does numerator simplify to ?

OpenStudy (anonymous):

you get 1(sqt 1-s) - 1+s ...right?

hartnn (hartnn):

the numerator now has \((1-\sqrt{s+1})(1+\sqrt{(s+1)})\) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the denominater is (1-s)s right?

hartnn (hartnn):

the denominator will be \(s \sqrt{s+1}(1+\sqrt{s+1})\) didn't u get this ?

hartnn (hartnn):

in the numerator, you use (a+b)(a-b)= a^2-b^2

OpenStudy (anonymous):

hmm I'm lost

hartnn (hartnn):

focus on numerator first ?

OpenStudy (anonymous):

Just ot be sure that is what you started out with right? as the equation \[(1\div \sqrt{1+s}) - 1 \div s\]

hartnn (hartnn):

sorry, i got to go the numerator becomes 1- (s+1) = -s this 's' gets cancelled with denominator's 's' and then if you plug in s=0, you will not get 0/0 form. good luck :)

hartnn (hartnn):

yes, thats what i started with...

hartnn (hartnn):

= (1-sqrt (s+1)-1)/ [s sqrt(s+1)] will be next step

OpenStudy (anonymous):

hold up I think i got it, what did you end up with your final answer?

OpenStudy (dumbcow):

limit is -1/2

OpenStudy (anonymous):

err I got 0

OpenStudy (dumbcow):

well following @hartnn steps you should end up with: \[\frac{-s}{s \sqrt{1+s}(1+\sqrt{1+s})}\] the "s" cancels plug in s=0